# A Survey on Deep Learning Hardware Accelerators for Heterogeneous HPC Platforms

### v2

@@ 4.4. Accelerators based on open-hardware RISC-V
rounded/.style=ellipse,draw, squared/.style=rectangle,draw, qtree, [RISC-V for Deep Learning [Specialized ISA & SIMD [ Dustin (ottaviDustin16CoresParallel2023, ), Marsellus (conti22124TOPS2023, )
Kraken (dimauroKrakenDirectEvent2022, ), Manticore (zarubaManticore4096CoreRISCV2021, )
Celerity (davidsonCelerityOpenSource511Core2018, ), Tenstorrent (vasiljevicComputeSubstrateSoftware2021, )
exSDOTP (bertacciniMiniFloatNNExSdotpISA2022, ), Esperanto (ditzelAcceleratingMLRecommendation2022, )
CNC (chenEightCoreRISCVProcessor2022, ), Cococcioni et al. (cococcioniLightweightPositPorcessing2021, )
PERCIVAL (mallasenPERCIVAL2022, ), Wang et al. (wangWinogradBasedConvolution2021, )
Tang et al. (tangGeneralPurposeGraphConvolutionNeuralAccelerator2022, ), RISC-VTF (JiaoRISCVExtensionTransformer2021, )
Paulin et al. (paulinRNNBasedRadioResource2021, ), Occamy (paulinOccamy432Core282024, )
Zhou et al. (zhouRISCVBasedFullyParallel2023, ), speedAI240 (snelgroveSpeedAI2402Petaflop30Teraflops2023, )
DECADES (gaoDECADES67mm246TOPS2023a, ), CIFER (liCIFERCacheCoherent12nm2023, )
RDCIM (yiRDCIMRISCVSupported2024d, ), Shaheen (valenteHeterogeneousRISCVBased2024, )
FlexNN (nadaliniTOPSRISCVParallel2023a, )
] ], [Vector Co-Processor [ Lee et al. (lee201445nm, ), AVA (lazoAdaptableRegisterFile2022, )
@@
Ara (cavalcanteAra1GHzScalable2020, ), Perotti et al. (perottiNewAraVector2022, )
@@
Garofalo et al. (garofaloHeterogeneousInMemoryComputing2022, ), Vega (rossiVegaTenCoreSoC2022, )
@@
Bruschi et al. (bruschiEndtoEndDNNInference2022, ), GAP9 (gap9Product, )
@@
Simba (shao_micro19, ), Lee et al. (lee64TOPSEnergyEfficientTensor2022, )
Tambe et al. (tambe2212nm182023, )
@@
] ], [L3-Coupled [ Gonzalez et al. (gonzalez16mm106GOPS2021, )
@@
Dos Santos et al. (dossantos1412nmLinuxSMPCapable2024, )
@@
Figure 5 reports a synthetic taxonomy of the RISC-V-based architectures for DL acceleration discussed in this survey.
@@ 4.4.1. RISC-V ISA extensions for (Deep) Learning
(Deep) neural networks are often limited by the computing and memory resources used by their large number of weights. Weight compression via alternative or quantized number representations is often adopted to speed up and optimize the performance of neural network models. Works in (cococcioniLightweightPositPorcessing2021, ; mallasenPERCIVAL2022, ) propose ISA extensions for posit numbers which can be used to do weight compression. Posit is an alternative representation to the standard IEEE floating point format for real numbers. Posit numbers need fewer bits to obtain the same precision or dynamic range of IEEE floats allowing them to store more weights in a same-sized memory. For example, the work in (cococcioniLightweightPositPorcessing2021, ) provides an efficient conversion between 8- or 16-bit posits and 32-bit IEEE floats or fixed point formats with little loss in precision leading to a 10x speedup in inference time. Other works directly address the compute-intensive parts of different neural networks, in particular CNNs, GCNs, and transformers. (wangWinogradBasedConvolution2021, ) proposes a new Winograd-based convolution instruction to speed up the time-consuming convolutional layers in CNNs. The matrix convolution between the CNN kernel and the input data cannot be executed efficiently using the standard RISC-V instructions. The proposed extension enables to compute a convolution producing a $2\times 2$ output using a $3\times 3$ kernel on a $4\times 4$ input in a single instruction using 19 clock cycles instead of multiple instructions totaling 140 cycles using the standard RISC-V ISA. (tangGeneralPurposeGraphConvolutionNeuralAccelerator2022, ) addresses the computational bottlenecks of GCNs by designing a set of general-purpose instructions for GCNs that mitigate the compute inefficiencies in aggregating and combining feature vectors. As such the authors combine the software programmability given by the RISC-V ISA with the compute efficiency of GCN accelerators. Similarly, (JiaoRISCVExtensionTransformer2021, ) focuses on transformer models. Notably, the extension comprises instructions to accelerate the well-known ReLU activation and softmax functions. Paulin et al. (paulinRNNBasedRadioResource2021, ) performs a similar task but focuses on RNNs.
@@
Following the trend to use aggressive quantization schemes to accelerate inference of DNNs, many ISA extensions focus on low-bit-width arithmetics to implement Quantized Neural Networks (QNNs), often combined with multi-core parallel execution to further boost performance and efficiency. Several developments augment the PULP RI5CY core used in Vega (rossiVegaTenCoreSoC2022, ) to improve its energy efficiency on QNNs. Marsellus (conti22124TOPS2023, ) (16 cores) and Kraken (dimauroKrakenDirectEvent2022, ) (8 cores) use Xpulpnn, an ISA extension for low-bitwidth (2/4-bit) integer dot-products used to accelerate symmetric precision QNNs, which is further extended in FlexNN (nadaliniTOPSRISCVParallel2023a, ). Dustin (ottaviDustin16CoresParallel2023, ) (16 cores) also exploits a similar concept, but it also introduces a lockstep mechanism to operate all the cores in a SIMD fashion, further increasing their efficiency. Shaheen (valenteHeterogeneousRISCVBased2024, ) exploits the same techniques in a more powerful SoC dedicated to applications for unmanned aerial vehicles. Many architectures exploit ISA extensions for faster RISC-V based DL workload execution in the context of many-core architectures where a large number of cores cooperate: academic examples include Manticore (zarubaManticore4096CoreRISCV2021, ), Occamy (paulinOccamy432Core282024, ), which includes tens-to-hundreds of simple 32-bit RISC-V cores, and CIFER (liCIFERCacheCoherent12nm2023, ), and DECADES (gaoDECADES67mm246TOPS2023a, ) which focus on exploiting tens of more complex 64-bit RISC-V cores. On the other hand, this approach is also popular with emerging commercial platforms, such as Celerity (davidsonCelerityOpenSource511Core2018, ), Esperanto (ditzelAcceleratingMLRecommendation2022, ), Tenstorrent (vasiljevicComputeSubstrateSoftware2021, ), and speedAI240 (snelgroveSpeedAI2402Petaflop30Teraflops2023, ). All these architectures are targeted at datacenter-based training and batch inference of large DNNs, Transformers, and Large Language Models (LLMs): hence, they typically focus on floating point multiply-accumulate and dot-product operations, such as exSDOTP (bertacciniMiniFloatNNExSdotpISA2022, ). Finally, a growing trend is to integrate digital in-memory computing (IMC) devices inside the pipeline of RISC-V processors as instruction set extensions. Two notable early examples at the silicon prototype maturity level are given by RDCIM (yiRDCIMRISCVSupported2024d, ) and Zhou et al. (zhouRISCVBasedFullyParallel2023, ).
@@ 4.4.2. RISC-V Vector Co-processors
Vector co-processors represent a sort of natural architectural target for Deep Learning-oriented RISC-V acceleration. AVA (lazoAdaptableRegisterFile2022, ), Vitruvius+ (minerviniVitruviusAreaEfficientRISCV2023, ), Ara (cavalcanteAra1GHzScalable2020, ; perottiNewAraVector2022, ; perottiYunOpenSource64Bit2023, ), Xvpfloat (guthmullerXvpfloatRISCVISA2024, ), MANIC (gobieskiMANIC19Mu2023, ) are academic vector co-processors meant to accelerate the full RISC-V V extension for vectorizable applications, including DL. Commercial RISC-V vector processors mainly targeted at HPC markets, such as Xuantie (chenXuantie910CommercialMultiCore2020, ), and Ventana (ventanaProduct, ), have recently started appearing. In addition, vector co-processors explicitly tailored for DL, like Spatz (cavalcanteSpatzCompactVector2022, ) and Arrow (assirArrowRISCVVector2021, ), have been developed. The former, in particular, focuses only on a subset of the V extension and on 32-bit data, capturing better opportunities for energy efficiency. Further pushing this trend, Vecim (wang30Vecim2892024, ) uses an IMC array to implement part of a reduced-precision (FP16) DL-dedicated vector extension for RISC-V.
@@ 4.4.3. RISC-V Memory-coupled Neural Processing Units (NPUs)
Concerning the tightest kind of memory coupling, at L1, most proposals in the state-of-the-art are based on the Parallel Ultra-Low Power (PULP) template, and devote significant effort to enabling fast communication between RISC-V cores and accelerators. Vega (rossiVegaTenCoreSoC2022, ) is a prototype system coupling 9 RI5CY cores with a quantized DNN convolutional NPU sharing directly L1 scratchpad memory with the cores. GreenWaves Technologies GAP9 (gap9Product, ) is a commercial product, targeted at the wearables market, that follows the same line with many architectural improvements and a redesigned AI NPU for QNNs – leading to the product achieving the best performance and energy efficiency in the public TinyMLPerf Tiny 1.0 challenge as of this writing (tinyMLPerf2023, ). Archimedes (prasadSpecializationMeetsFlexibility, ) and Siracusa (prasadSiracusa16Nm2024, ) propose a large AI hardware accelerator for performance-hungry Extended Reality applications, directly coupled with an embedded non-volatile memory (MRAM) in an “at-memory” configuration. RedMulE (tortorellaRedMulECompactFP162022a, ; tortorellaRedMuleMixedPrecisionMatrixMatrix2023, ), integrated into the Darkside prototype (garofaloDARKSIDEHeterogeneousRISCV2022, ), follows the same principles but focusing on floating-point computation to support training as well as inference. Garofalo et al. (garofaloHeterogeneousInMemoryComputing2022, ) and Bruschi et al. (bruschiEndtoEndDNNInference2022, ) integrate in-memory-computing PCM-based NPUs, the latter simulating a system scaled up to match the size of server-class hardware.
@@
Figure 6. Performance and power consumption of prototypes of several State-of-the-Art Deep Learning acceleration architectures discussed in Section 4.4.
@@
Moving the shared memory from L1 to L2/L3, there are other NPU solutions in the state-of-the-art. In these architectures, RISC-V open-source cores are flexible and easy to be integrated. Systems exploiting this template include, for example, SNCPU (juSystolicNeuralCPU2023, ), which builds a hybrid system that can act as either a set of 10 RISC-V cores or be reconfigured in a systolic NPU. Gonzalez et al. (gonzalez16mm106GOPS2021, ) and Genc et al. (gencGemminiEnablingSystematic2021, ) exploit a systolic array generation tool, Gemmini, to generate systolic arrays to accelerate DNNs and coupled with a RISC-V core by exploiting a shared L3 or L2 memory, respectively. Simba (shao_micro19, ) follows a similar template and is also meant to be scaled towards server-grade performance using the integration of chiplets on multi-chip modules. ESP (giriESP4MLPlatformBasedDesign2020, ; giriAcceleratorIntegrationOpenSource2021, ) and Tambe et al. (tambe2212nm182023, ) also focus on integrating hardware accelerators and NPUs in large-scale Network-on-Chips using RISC-V cores as computing engines. Axelera AI propose a so-far unique architecture in the spectrum of RISC-V couple NPUs: a L2 shared-memory accelerator exploiting digital SRAM-based IMC, called Metis AIPU (hager11MetisAIPU2024, ), targeting high-performance edge and server-based inference. On the other end of the spectrum, SamurAI (miro-panadesSamurAIVersatileIoT2022, ), TinyVers (jainTinyVersTinyVersatile2023, ), and DIANA (houshmandDIANAEndtoEndHybrid2023, ) build up AI-IoT systems composed of a microcontroller and L2-coupled NPUs (in the case of DIANA both an analog SRAM-IMC-based accelerator and a conventional digital one). Kraken (dimauroKrakenDirectEvent2022, ) couples the aforementioned RISC-V ISA-extended cluster with specialized L2-coupled Spiking Neural Network (SNN) and Ternary Neural Network (TNN) accelerators.
@@ 4.4.4. Summary
Figure 6 summarizes performance, power, technological maturity, and data types used in the architectures discussed in this section. RISC-V-based solutions occupy essentially the full spectrum of DL architectures ranging from 10 mW microcontrollers up to 100 W SoCs meant to be integrated as part of HPC systems. So far, most of the research has focused on the lower end of this spectrum, striving for the best energy efficiency. We can observe how efficiency is strongly correlated with architectural techniques yielding accuracy (e.g., data bit-width reduction & quantization) and with the usage of emerging computational paradigms such as in-memory computing. Table 5 summarizes all quantitative information available on the discussed architectures; whenever multiple operating points have been reported for a single architecture, the table reports always the highest performance (for consistency, this applies to both performance and energy efficiency values).
5. Acceleration-based on Emerging Technologies
To design efficient DNN hardware accelerators, combining optimized memory architectures and processing modules is crucial to achieve high speed at reasonable costs and power dissipation. Such architectures must be designed taking into account the large amount of memory necessary to store the input feature maps, weights, and intermediate results generated by the convolutional layers in a DNN. Moreover, managing DNN computational models causes a large number of data movements between the memory and the processing elements, often posing a challenge in terms of achievable speed performance, energy consumption, and memory bandwidth. For these reasons, several innovative memory architectures and technologies have recently emerged to increase memory capacity and data bandwidth, to reduce memory access latency, and potentially to improve the power efficiency. In this Section, we discuss two efficient technologies: Processing-in-Memory and In-Memory Computing.

### v3

@@ 4.4. Accelerators based on open-hardware RISC-V
rounded/.style=ellipse,draw, squared/.style=rectangle,draw, qtree, [RISC-V for Deep Learning [Specialized ISA & SIMD [ Dustin (ottaviDustin16CoresParallel2023, ), Marsellus (conti22124TOPS2023, ), Kraken (dimauroKrakenDirectEvent2022, )
Manticore (zarubaManticore4096CoreRISCV2021, ), Celerity (davidsonCelerityOpenSource511Core2018, ), Tenstorrent (vasiljevicComputeSubstrateSoftware2021, )
exSDOTP (bertacciniMiniFloatNNExSdotpISA2022, ), Esperanto (ditzelAcceleratingMLRecommendation2022, ), CNC (chenEightCoreRISCVProcessor2022, )
Cococcioni (cococcioniLightweightPositPorcessing2021, ), PERCIVAL (mallasenPERCIVAL2022, ), Wang (wangWinogradBasedConvolution2021, )
Tang (tangGeneralPurposeGraphConvolutionNeuralAccelerator2022, ), RISC-VTF (JiaoRISCVExtensionTransformer2021, ), Paulin (paulinRNNBasedRadioResource2021, )
Occamy (paulinOccamy432Core282024, ), Zhou (zhouRISCVBasedFullyParallel2023, ), speedAI240 (snelgroveSpeedAI2402Petaflop30Teraflops2023, )
DECADES (gaoDECADES67mm246TOPS2023a, ), CIFER (liCIFERCacheCoherent12nm2023, ), RDCIM (yiRDCIMRISCVSupported2024d, )
Shaheen (valenteHeterogeneousRISCVBased2024, ), FlexNN (nadaliniTOPSRISCVParallel2023a, )
] ], [Vector Co-Processor [ Lee (lee201445nm, ), AVA (lazoAdaptableRegisterFile2022, )
@@
Ara (cavalcanteAra1GHzScalable2020, ), Perotti (perottiNewAraVector2022, )
@@
Garofalo (garofaloHeterogeneousInMemoryComputing2022, ), Vega (rossiVegaTenCoreSoC2022, )
@@
Bruschi (bruschiEndtoEndDNNInference2022, ), GAP9 (gap9Product, )
@@
Simba (shao_micro19, ), Lee (lee64TOPSEnergyEfficientTensor2022, )
Tambe (tambe2212nm182023, )
@@
] ], [L3-Coupled [ Gonzalez (gonzalez16mm106GOPS2021, )
@@
Dos Santos (dossantos1412nmLinuxSMPCapable2024, )
@@
Figure 4 reports a synthetic taxonomy of representative RISC-V-based accelerators for DL.
@@ 4.4.1. RISC-V ISA extensions for (Deep) Learning
Works in (cococcioniLightweightPositPorcessing2021, ; mallasenPERCIVAL2022, ) propose ISA extensions for posit numbers which can be used to do weight compression. Posit numbers need fewer bits to obtain the same precision or dynamic range of IEEE floats allowing them to store more weights in a same-sized memory. For example, the work in (cococcioniLightweightPositPorcessing2021, ) provides an efficient conversion between 8- or 16-bit posits and 32-bit IEEE floats or fixed point formats with little loss in precision leading to a 10x speedup in inference time. Other works directly address the compute-intensive parts of different neural networks, in particular CNNs, GCNs, and transformers. The new Winograd-based convolution instruction proposed in (wangWinogradBasedConvolution2021, ) enables to compute a convolution producing a $2\times 2$ output using a $3\times 3$ kernel on a $4\times 4$ input in a single instruction using 19 clock cycles instead of multiple instructions totaling 140 cycles using the standard RISC-V ISA. The set of general-purpose instructions for GCNs designed in (tangGeneralPurposeGraphConvolutionNeuralAccelerator2022, ) mitigate the compute inefficiencies in aggregating and combining feature vectors. As such the authors combine the software programmability given by the RISC-V ISA with the compute efficiency of GCN accelerators. Similarly, (JiaoRISCVExtensionTransformer2021, ) focuses on transformer models. Notably, the extension comprises instructions to accelerate the well-known ReLU activation and softmax functions. Paulin et al. (paulinRNNBasedRadioResource2021, ) performs a similar task but focuses on RNNs.
@@
Many ISA extensions focus on low-bit-width arithmetics to accelerate inference of Quantized Neural Networks (QNNs), often combined with multi-core parallel execution to further boost performance and efficiency. Several developments augment the PULP RI5CY core used in Vega (rossiVegaTenCoreSoC2022, ) to improve its energy efficiency on QNNs. Marsellus (conti22124TOPS2023, ) (16 cores) and Kraken (dimauroKrakenDirectEvent2022, ) (8 cores) use Xpulpnn, an ISA extension for low-bitwidth (2/4-bit) integer dot-products used to accelerate symmetric precision QNNs, which is further extended in FlexNN (nadaliniTOPSRISCVParallel2023a, ). Dustin (ottaviDustin16CoresParallel2023, ) (16 cores) also exploits a similar concept, but it also introduces a lockstep mechanism to operate all the cores in a SIMD fashion, further increasing their efficiency. Shaheen (valenteHeterogeneousRISCVBased2024, ) exploits the same techniques in a more powerful SoC dedicated to applications for unmanned aerial vehicles. Many architectures, like Manticore (zarubaManticore4096CoreRISCV2021, ), Occamy (paulinOccamy432Core282024, ), CIFER (liCIFERCacheCoherent12nm2023, ), and DECADES (gaoDECADES67mm246TOPS2023a, ), exploit ISA extensions for faster RISC-V based DL workload execution in the context of many-core architectures where a large number of cores cooperate On the other hand, this approach is also popular with emerging commercial platforms, such as Celerity (davidsonCelerityOpenSource511Core2018, ), Esperanto (ditzelAcceleratingMLRecommendation2022, ), Tenstorrent (vasiljevicComputeSubstrateSoftware2021, ), and speedAI240 (snelgroveSpeedAI2402Petaflop30Teraflops2023, ). All these architectures are targeted at datacenter-based training and batch inference of large DNNs, Transformers, and Large Language Models (LLMs): hence, they typically focus on floating point multiply-accumulate and dot-product operations, such as exSDOTP (bertacciniMiniFloatNNExSdotpISA2022, ). Finally, a growing trend is to integrate digital in-memory computing (IMC) devices inside the pipeline of RISC-V processors as instruction set extensions. Two notable early examples at the silicon prototype maturity level are given by RDCIM (yiRDCIMRISCVSupported2024d, ) and Zhou et al. (zhouRISCVBasedFullyParallel2023, ).
@@ 4.4.2. RISC-V Vector Co-processors
Vector co-processors represent a sort of natural architectural target for DL-oriented RISC-V acceleration. AVA (lazoAdaptableRegisterFile2022, ), Vitruvius+ (minerviniVitruviusAreaEfficientRISCV2023, ), Ara (cavalcanteAra1GHzScalable2020, ; perottiNewAraVector2022, ; perottiYunOpenSource64Bit2023, ), Xvpfloat (guthmullerXvpfloatRISCVISA2024, ), MANIC (gobieskiMANIC19Mu2023, ) are academic vector co-processors meant to accelerate the full RISC-V V extension for vectorizable applications, including DL. Commercial RISC-V vector processors mainly targeted at HPC markets, such as Xuantie (chenXuantie910CommercialMultiCore2020, ), and Ventana (ventanaProduct, ), have recently started appearing. In addition, vector co-processors explicitly tailored for DL, like Spatz (cavalcanteSpatzCompactVector2022, ) and Arrow (assirArrowRISCVVector2021, ), have been developed. The former, in particular, focuses only on a subset of the V extension and on 32-bit data, capturing better opportunities for energy efficiency. Further pushing this trend, Vecim (wang30Vecim2892024, ) uses an IMC array to implement part of a reduced-precision (FP16) DL-dedicated vector extension for RISC-V.
@@ 4.4.3. RISC-V Memory-coupled Neural Processing Units (NPUs)
Concerning the tightest kind of memory coupling, at L1, most proposals in the state-of-the-art are based on the Parallel Ultra-Low Power (PULP) template, and devote significant effort to enabling fast communication between RISC-V cores and accelerators. Representative system architectures designed in this way are available at several levels of maturity, like the prototypes Vega (rossiVegaTenCoreSoC2022, ) and Darkside (garofaloDARKSIDEHeterogeneousRISCV2022, ), the commercial products GreenWaves Technologies GAP9 (gap9Product, ), Archimedes (prasadSpecializationMeetsFlexibility, ) and Siracusa (prasadSiracusa16Nm2024, ), and the simulation templates  (garofaloHeterogeneousInMemoryComputing2022, ) and  (bruschiEndtoEndDNNInference2022, ).
@@
Figure 5. Performance and power consumption of SotA DL accelerators based on open-HW RISC-V.
@@
Moving the shared memory from L1 to L2/L3, there are other NPU solutions in the state-of-the-art. For example, SNCPU (juSystolicNeuralCPU2023, ), can act as either a set of 10 RISC-V cores or be reconfigured in a systolic NPU. In  (gonzalez16mm106GOPS2021, ) and  (gencGemminiEnablingSystematic2021, ), systolic arrays generated by Gemmini are coupled with a RISC-V core by exploiting a shared L3 or L2 memory, respectively. Simba (shao_micro19, ) is also meant to be scaled towards server-grade performance using the integration of chiplets on multi-chip modules. ESP (giriESP4MLPlatformBasedDesign2020, ; giriAcceleratorIntegrationOpenSource2021, ) and  (tambe2212nm182023, ) also focus on integrating hardware accelerators and NPUs in large-scale Network-on-Chips using RISC-V cores as computing engines. Axelera AI propose a so-far unique architecture that uses a L2 shared-memory accelerator exploiting digital SRAM-based IMC, called Metis AIPU (hager11MetisAIPU2024, ). SamurAI (miro-panadesSamurAIVersatileIoT2022, ), TinyVers (jainTinyVersTinyVersatile2023, ), and DIANA (houshmandDIANAEndtoEndHybrid2023, ) build up AI-IoT systems composed of a microcontroller and L2-coupled NPUs. Kraken (dimauroKrakenDirectEvent2022, ) couples the RISC-V ISA-extended cluster with specialized L2-coupled Spiking Neural Network (SNN) and Ternary Neural Network (TNN) accelerators.
@@ 4.4.4. Summary
Figure 5 clearly shows that RISC-V-based solutions occupy essentially the full spectrum of DL architectures ranging from 10 mW microcontrollers up to 100 W SoCs meant to be integrated as part of HPC systems. So far, most of the research has focused on the lower end of this spectrum, striving for the best energy efficiency. We can observe how efficiency is strongly correlated with architectural techniques yielding accuracy (e.g., data bit-width reduction & quantization) and with the usage of emerging computational paradigms such as in-memory computing. Table 5 compares the above discussed architectures quantitatively and reports their highest performance and energy efficiency values.
5. Accelerators based on Emerging Technologies
To design efficient DNN hardware accelerators, combining optimized memory architectures and processing modules is crucial to achieve high speed at reasonable costs and power dissipation. Such architectures must be designed taking into account the large amount of memory necessary to store the input feature maps, weights, and intermediate results generated by the convolutional layers in a DNN. Moreover, managing DNN computational models causes a large number of data movements between the memory and the processing elements, often posing a challenge in terms of achievable speed performance, energy consumption, and memory bandwidth. For these reasons, several innovative memory architectures and technologies have recently emerged to increase memory capacity and data bandwidth, to reduce memory access latency, and potentially to improve the power efficiency. In this Section, we discuss several technologies: Processing-in-Memory and In-Memory Computing (see Fig. 6), Neuromorphic accelerators, approaches based on Multi-Chip Modules, and Quantum and Photonic computing.
{forest}
SIMDRAM(Hajinazar_2021, ), PIM-DRAM(Roy_2021, )
] ] [Logic die (3D-DRAM) [ Neurocube(Kim2016, ), Tetris(Gao2017, )
NeuralHMC(Min2019, ), VIMA(Cordeiro2021, )
] ] [Bank (3D-DRAM) [ Newton(He2020, )
HBM-PIM(Kwon2021, )
] ] ], [IMC [SRAM [Orlando (desoli17isscc, ),
STM-IMC (desoli2023isscc, ) ] ] [PCM [ Hermes (khaddam2022hermes, ),
NeuralPIM (cao2022tcomp, ), NeuRRAM (wan2022nature, )